OpenStudy (anonymous):

Solve: -2 1/2w=3 1/3

OpenStudy (anonymous):

2 1/2 is the same as two and a half which is the same as 5/2 3 1/3 is the same as three and one-third which is the same as 10/3

OpenStudy (anonymous):

so its 10/3

OpenStudy (anonymous):

?

OpenStudy (anonymous):

nope. -5/2 w = 10/3 so w = ?

OpenStudy (anonymous):

idk i hate math!

OpenStudy (anonymous):

okay, cross multiply. -w = 10*2/5*3 = 20/15 or w = -20/15 w = -4/3